### v1

@@ Abstract
[Abstract]Background: The integration of haptic feedback into robot-assisted minimally invasive surgery (RAMIS) has long been limited by challenges in accurately rendering forces and ensuring system safety. The need for robust, high-fidelity haptic systems is critical for enhancing the precision and reliability of teleoperated surgical tools.
Methods: In this study, we present a Nonlinear Impedance Matching Approach (NIMA) designed to improve force rendering by accurately modelling complex tool-tissue interactions. Based on our previously validated Impedance Matching Approach (IMA), our novel NIMA method includes nonlinear dynamics to capture and render tool-tissue forces effectively.
Results: NIMA improves force feedback accuracy with a mean absolute error (MAE) of 0.01 $\pm$ 0.02 N, achieving a 95% reduction in MAE compared to IMA. Furthermore, NIMA effectively eliminates haptic "kickback" by ensuring no force is applied by the haptic device to the user’s hand when they release the handle, enhancing both patient safety and user comfort.
Conclusion: NIMA’s ability to account for nonlinearities in tool-tissue interactions provides an improvement in force fidelity, responsiveness, and precision across various surgical conditions. Our findings promote the advancement of haptic feedback systems for robotic surgery, offering a realistic and reliable interface for robot-assisted surgical procedures.
@@
Haptics, Impedance Matching Approach, Force Rendering, Human-in-the-loop, Surgical Robotics, Minimally Invasive Surgery, Laparoscopy.
articletype: Original Articlejournal: The International Journal of Medical Robotics and Computer Assisted Surgeryvolume: 00
@@ 1 Introduction
Advances in haptic technology are central to enhancing patient safety, surgical precision, and operator performance in robot-assisted minimally invasive surgery (RAMIS) colan2024tactile, Amirabdollahian2017. The integration of haptic or force feedback into teleoperated RAMIS platforms provides substantial clinical and scientific advantages, including real-time assistance to surgeons, improved perception of tissue consistency, and automatic acquisition of tissue mechanical properties Okamura2004, torkaman2023embedded, roshanfar2025learning. Several studies have shown that haptic feedback enhances the consistency, accuracy, and efficiency of tasks such as knot tying, thereby reducing the likelihood of inadvertent tissue damage Mei2011. Moreover, it improves overall surgical outcomes by enabling finer motor control and augmenting visual feedback, ultimately reducing patient risk hooshiar2019haptic, Meijden2009, Zhou2020. Despite these benefits, most current teleoperated surgical systems still operate without direct haptic feedback due to challenges in maintaining closed-loop stability and ensuring patient safety Pacchierotti2016. For instance, while next-generation RAMIS platforms such as the da Vinci 5 system have begun integrating limited force-sensing capabilities for enhanced situational awareness, they do not yet provide continuous force reflection to the operator. Consequently, the integration of robust and safe haptic feedback remains one of the most critical technological challenges in achieving full sensory immersion and automation in robotic surgery Meijden2009, Gumbs2021.
The regulatory approval of haptic systems as human-in-the-loop components in RAMIS continues to face significant challenges, primarily due to the difficulty of accounting for variations in human dynamics during validation and testing Pacchierotti2017, Okamura2004. Stability concerns in force-feedback teleoperation and uncertainties in force estimation further complicate the safe deployment of such systems in clinical environments Bahar2020. Designing haptic devices that simultaneously ensure stability, transparency, and intuitive interaction while satisfying the diverse ergonomic and performance requirements of surgeons remains a major engineering challenge Selim2023. Moreover, variations in grip strength, reaction time, and motion control among individual users introduce additional uncertainties that hinder regulatory standardization and certification Chae2018, Fu2012. The persistent absence of high-fidelity haptic sensation in most commercial RAMIS platforms continues to limit their ability to fully replicate the tactile experience of open surgery Lai2022. As the adoption of robotic surgery expands globally, the integration of safe, robust, and validated haptic feedback systems (HFS) into next-generation surgical robots has become increasingly imperative to bridge this sensory gap and enhance both training and operative performance Abiri2019.
@@ 1.1 Related Studies
In the past decade, advances in commercial platforms and AI-enabled sensing methods have significantly advanced the provision of robust haptic feedback in RAMIS. Recent reviews and meta-analyses consistently show that augmenting vision with force or tactile information improves task performance, reduces peak forces, and can reduce errors in simulated and pre-clinical settings patel2022haptic, bergholz2023benefits, colan2024tactile, boul2025role, dagnino2024robot. In Table 1 we have synthesized a non-exhaustive list of current main approaches, i.e., direct tip sensing, proximal or joint-level sensing, model-based sensor-free estimation, vision-based force estimation, and purely visual or pseudo-haptic cueing, and compared them in terms of algorithmic complexity, hardware-agnostic deployment, sensor dependence, and whether they rely on proximal or distal measurements relative to the tool–tissue interface.
@@
Instrument-integrated force sensing represents the most intuitive route to robust haptic feedback, and several recent studies have quantified its impact in realistic surgical tasks. Awad et al. evaluated a new generation of robotic instruments with tip-mounted force sensors and demonstrated that force-feedback substantially reduced mean and maximal forces during retraction, dissection, and suturing on inanimate and ex-vivo models across experience levels awad2024evaluation. Systematic reviews of tactile and kinesthetic feedback in RAMIS confirm that such “true” haptics can improve precision and reduce error, but they also highlight integration complexity, sterilization constraints, and platform-specific hardware as major barriers to widespread adoption colan2024tactile, boul2025role. These solutions offer high-fidelity distal sensing at the cost of high mechanical and electronic complexity and low hardware agnosticity. To avoid modifying the distal instrument, several groups have pursued proximal or sensorless approaches that reconstruct interaction forces from joint torques, actuator currents, or high-level motion descriptors. Pisla et al. proposed an AI-based sensorless force-feedback strategy for robot-assisted esophagectomy that infers interaction forces from robot-side signals and uses them to generate haptic cues without any additional sensors on the surgical tools pisla2025ai. Yan et al. developed a deep-learning model based on an ISSA-optimized backpropagation Neural Network (NN) to predict clamp forces on soft tissue using clamp motion parameters (contact area, speed, displacement, time) measured during compression tests yan2025robust. These methods, summarized under “sensorless/model-based” in Table 1, tend to be more hardware-agnostic at the tool level and easier to retrofit to existing robots, but they require accurate dynamic models or extensive training data, careful calibration, and explicit handling of tissue variability to remain robust in vivo.
@@
A complementary trend is the use of vision-based force estimation, where endoscopic images (sometimes combined with robot kinematics) are used to infer contact forces or detect over-force events. Early work by Chua et al. combined vision and robot state to learn a mapping to interaction forces in RAMIS chua2021toward, and subsequent studies have refined these ideas. Masui et al. showed that deep learning models trained on laparoscopic images of a porcine kidney can estimate manipulation forces and detect over-force events when the region of interest is restricted around the tool tips masui2024vision. Yang et al. introduced a contact-conditional framework that uses vision with local stiffness models and optionally joint torque information to predict forces in minimally invasive telesurgery, achieving high accuracy on contact detection and sub-10% force prediction errors yang2024vision. Ding et al. surveyed this emerging field and emphasized the potential of vision-based contact force detection to reduce hardware complexity and enable platform-agnostic haptic augmentation ding2025vision. In Table 1, these approaches are characterized by high algorithmic complexity but minimal additional hardware, with distal sensing achieved indirectly through image analysis.
@@
Vision-based methods naturally connect to pseudo-haptic and visual-only feedback strategies, where the goal is not to reconstruct forces precisely but to shape the surgeon’s perception of tissue stiffness or danger through visual deformation cues or other modalities. Masui et al. explicitly frame their vision-based force estimation as an attempt to make surgeons’ existing “pseudo-haptic” impressions more explicit and actionable masui2024vision. Trute et al. studied how soft-tissue visual cues in minimal-invasive and robotic surgery can support haptic perception and inform the design of visual augmentations trute2024visual. Broader reviews of haptic technologies for surgical simulation and robotics also stress that pseudo-haptic and multi-modal sensory substitution (e.g., vibrotactile cues) can reduce hardware demands while still improving training and performance boul2025role. These approaches are highly hardware-agnostic and low-cost but rely heavily on robust computer vision and careful human factor design rather than physically accurate force reconstruction.
@@
Distal force sensing awad2024evaluation, colan2024tactile, bergholz2023benefits High (H) Low (L) High (H) Distal (D)
Proximal/Joint-space sensing pisla2025ai, yan2025robust, patel2022haptic Med (M) Med (M) Robot-side only (R) Proximal (P)
@@
Vision-based force estimation chua2021toward, masui2024vision, yang2024vision, ding2025vision High (H) High (H) Low physical (L) + High data (Hd) Distal (Dind)
Pseudo-haptics / visual cues masui2024vision, trute2024visual, boul2025role Low–Med (L–M) Very High (VH) Low (L) Visual distal (V)
Hybrid multi-modal fusion patel2022haptic, ding2025vision, boul2025role, dagnino2024robot High (H) Med (M) Mixed (Mx) P + D
@@
Taken together, recent literature suggests that there is no single “best” method for robust haptic feedback, instead, different clinical, regulatory, and engineering constraints naturally map to different regions of the design space. For new robotic platforms where instrument redesign is feasible and the cost per-unit is acceptable, instrument-integrated force sensors provide the most direct, distal measurements and the strongest evidence of performance benefits awad2024evaluation, colan2024tactile, bergholz2023benefits. For retrofitting legacy systems or ensuring cross-platform deployability, proximal sensing and sensorless learning-based models offer a pragmatic compromise between hardware-agnostic deployment and robustness, especially when trained on data that span the expected range of tasks and tissues pisla2025ai, yan2025robust, patel2022haptic. Vision-based and pseudo-haptic methods, finally, are attractive when hardware changes are impossible (e.g., closed commercial systems) or when force information is primarily needed for training, monitoring, or safety alerts rather than for continuous kinesthetic feedback masui2024vision, yang2024vision, ding2025vision, trute2024visual.
@@
More recently, Impedance Matching Approach (IMA) for force feedback has emerged as an effective strategy for achieving stable and transparent haptic interaction in RAMIS. It operates as a form of sensory substitution, enabling indirect force rendering by bypassing actuators on the leader side and excluding human components from the control loop Yin2018, Munawar2016, Prattichizzo2012. Unlike conventional direct force feedback methods, where force error directly commands actuator output, IMA employs closed-loop position control, converting the force error into a corresponding position error that drives the robot golahmadi2021tool, haouchine2018vision, Wilfinger1994. This formulation allows precise regulation of tool–tissue contact forces while maintaining system stability through an inner position-control loop Siciliano1996. Previous studies have validated the IMA framework through simulation and experimental evaluations, demonstrating its potential for stable and accurate haptic rendering Parsi2023. Complementary work has explored integrating tactile object recognition and purposeful haptic exploration to extract object features and improve force perception during teleoperation Pezzementi2011. Moreover, advanced haptic rendering algorithms have been shown to minimize feedback oscillations and enhance the fidelity of force cues, leading to more realistic and responsive interaction Liu2022. Building on this, the linear IMA method has recently demonstrated high precision and stability for single-axis force components in surgical teleoperation tasks sayadi2020impedance, motivating further development of nonlinear extensions for more complex tool–tissue interactions.
@@ 1.2 Contributions
This study introduces a generalized Nonlinear Impedance Matching Approach (NIMA), which extends the conventional IMA framework by incorporating nonlinear impedance parameters to more accurately capture the complex dynamics of instrument–tissue interactions encountered in RAMIS. As illustrated in Fig. 1, the proposed NIMA framework continuously identifies nonlinear impedance characteristics in real-time to enable precise rendering of multi-axis contact forces and stable bidirectional communication between the leader and follower modules. The framework was implemented and experimentally validated using commercially available laparoscopic tools, confirming its robustness in controlling three-dimensional contact forces and executing motion commands with high fidelity. The main contributions of this work are:
@@
    A novel NIMA that generalizes the traditional IMA framework to model nonlinear tool–tissue dynamics for improved force rendering in teleoperated surgery.
@@
    A complete leader–follower control architecture enabling real-time estimation and transmission of impedance parameters for stable and responsive haptic feedback.
@@
    Experimental validation demonstrating accurate 3D force control.
@@
Figure 1: Proposed Nonlinear Impedance Matching Approach (NIMA). The leader module (left) receives motion commands $\mathbf{X}$ from the operator and renders the estimated feedback force $\mathbf{f_{d}}=\mathbf{M}\mathbf{X}$, while the follower module (right) executes $\mathbf{X}$, measures $(\mathbf{f},\mathbf{X})$, and identifies nonlinear impedance parameters $\mathbf{M}$ in real time. This closed-loop structure enables stable, high-fidelity force feedback in robotic laparoscopy mazidi2024nonlinear.
2 Methodology

### v2

@@ Abstract
Integrating accurate haptic feedback into robot-assisted minimally invasive surgery (RAMIS) remains challenging due to difficulties in precise force rendering and ensuring system safety during teleoperation. We present a Nonlinear Impedance Matching Approach (NIMA) that extends our previously validated Impedance Matching Approach (IMA) by incorporating nonlinear dynamics to accurately model and render complex tool-tissue interactions in real-time. NIMA achieves a mean absolute error of 0.01 $\pm$ 0.02 N, representing a 95% reduction compared to IMA. Additionally, NIMA eliminates haptic “kickback” by ensuring zero force is applied to the user’s hand when they release the handle, enhancing both patient safety and operator comfort. By accounting for nonlinearities in tool-tissue interactions, NIMA significantly improves force fidelity, responsiveness, and precision across various surgical conditions, advancing haptic feedback systems for reliable robot-assisted surgical procedures.
@@
haptics , impedance matching approach , force rendering , human-in-the-loop , surgical robotics , minimally invasive surgery , laparoscopy
journal: Mechatronics
[…]
[aff1]organization=Surgical Robotics Laboratory (SRL), Mechanical Engineering Department, Concordia University, city=Montreal, state=QC, country=Canada
[…]
[aff2]organization=Surgical Performance Enhancement and Robotics (SuPER) Centre, Department of Surgery, McGill University, city=Montreal, state=QC, country=Canada
[…]
[aff3]organization=The Wilfred and Joyce Posluns Centre for Image Guided Innovation & Therapeutic Intervention (PCIGITI), The Hospital for Sick Children (SickKids), city=Toronto, state=ON, country=Canada
{highlights}
A novel Nonlinear Impedance Matching Approach (NIMA) for stable haptic feedback in robot-assisted surgery.
Neural network-based tool-tip force extraction isolating friction at the remote center of motion.
95% accuracy improvement over linear IMA with MAE of 0.01 N.
Elimination of haptic kickback ensuring operator safety and comfort.
@@ 1 Introduction
Advances in haptic technology are central to enhancing patient safety, surgical precision, and operator performance in robot-assisted minimally invasive surgery (RAMIS) [9, 2]. The integration of haptic or force feedback into teleoperated RAMIS platforms provides substantial clinical and scientific advantages, including real-time assistance to surgeons, improved perception of tissue consistency, and automatic acquisition of tissue mechanical properties [27, 40, 36]. Several studies have shown that haptic feedback enhances the consistency, accuracy, and efficiency of tasks such as knot tying, thereby reducing the likelihood of inadvertent tissue damage [24]. Moreover, it improves overall surgical outcomes by enabling finer motor control and augmenting visual feedback, ultimately reducing patient risk [18, 25, 46]. Despite these benefits, most current teleoperated surgical systems still operate without direct haptic feedback due to challenges in maintaining closed-loop stability and ensuring patient safety [28]. For instance, while next-generation RAMIS platforms such as the da Vinci 5 system have begun integrating limited force-sensing capabilities for enhanced situational awareness, they do not yet provide continuous force reflection to the operator. Consequently, the integration of robust and safe haptic feedback remains one of the most critical technological challenges in achieving full sensory immersion and automation in robotic surgery [25, 16].
The regulatory approval of haptic systems as human-in-the-loop components in RAMIS continues to face significant challenges, primarily due to the difficulty of accounting for variations in human dynamics during validation and testing [29, 27]. Stability concerns in force-feedback teleoperation and uncertainties in force estimation further complicate the safe deployment of such systems in clinical environments [4, 14]. Designing haptic devices that simultaneously ensure stability, transparency, and intuitive interaction while satisfying the diverse ergonomic and performance requirements of surgeons remains a major engineering challenge [38]. Moreover, variations in grip strength, reaction time, and motion control among individual users introduce additional uncertainties that hinder regulatory standardization and certification [7, 13]. The persistent absence of high-fidelity haptic sensation in most commercial RAMIS platforms continues to limit their ability to fully replicate the tactile experience of open surgery [20]. As the adoption of robotic surgery expands globally, the integration of safe, robust, and validated haptic feedback systems (HFS) into next-generation surgical robots has become increasingly imperative to bridge this sensory gap and enhance both training and operative performance [1].
@@ 1.1 Related Studies
In the past decade, advances in commercial platforms and AI-enabled sensing methods have significantly advanced the provision of robust haptic feedback in RAMIS. Recent reviews and meta-analyses consistently show that augmenting vision with force or tactile information improves task performance, reduces peak forces, and can reduce errors in simulated and pre-clinical settings [31, 5, 9, 6, 10]. In Table 1 we have synthesized a non-exhaustive list of current main approaches, i.e., direct tip sensing, proximal or joint-level sensing, model-based sensor-free estimation, vision-based force estimation, and purely visual or pseudo-haptic cueing, and compared them in terms of algorithmic complexity, hardware-agnostic deployment, sensor dependence, and whether they rely on proximal or distal measurements relative to the tool-tissue interface.
@@
Instrument-integrated force sensing represents the most intuitive route to robust haptic feedback, and several recent studies have quantified its impact in realistic surgical tasks. Awad et al. evaluated a new generation of robotic instruments with tip-mounted force sensors and demonstrated that force-feedback substantially reduced mean and maximal forces during retraction, dissection, and suturing on inanimate and ex-vivo models across experience levels [3]. Systematic reviews of tactile and kinesthetic feedback in RAMIS confirm that such “true” haptics can improve precision and reduce error, but they also highlight integration complexity, sterilization constraints, and platform-specific hardware as major barriers to widespread adoption [9, 6]. These solutions offer high-fidelity distal sensing at the cost of high mechanical and electronic complexity and low hardware agnosticity. To avoid modifying the distal instrument, several groups have pursued proximal or sensorless approaches that reconstruct interaction forces from joint torques, actuator currents, or high-level motion descriptors. Pisla et al. proposed an AI-based sensorless force-feedback strategy for robot-assisted esophagectomy that infers interaction forces from robot-side signals and uses them to generate haptic cues without any additional sensors on the surgical tools [34]. Yan et al. developed a deep-learning model based on an ISSA-optimized backpropagation Neural Network (NN) to predict clamp forces on soft tissue using clamp motion parameters (contact area, speed, displacement, time) measured during compression tests [43]. These methods, summarized under “sensorless/model-based” in Table 1, tend to be more hardware-agnostic at the tool level and easier to retrofit to existing robots, but they require accurate dynamic models or extensive training data, careful calibration, and explicit handling of tissue variability to remain robust in vivo.
@@
A complementary trend is the use of vision-based force estimation, where endoscopic images (sometimes combined with robot kinematics) are used to infer contact forces or detect over-force events. Early work by Chua et al. combined vision and robot state to learn a mapping to interaction forces in RAMIS [8], and subsequent studies have refined these ideas. Masui et al. showed that deep learning models trained on laparoscopic images of a porcine kidney can estimate manipulation forces and detect over-force events when the region of interest is restricted around the tool tips [22]. Yang et al. introduced a contact-conditional framework that uses vision with local stiffness models and optionally joint torque information to predict forces in minimally invasive telesurgery, achieving high accuracy on contact detection and sub-10% force prediction errors [44]. Ding et al. surveyed this emerging field and emphasized the potential of vision-based contact force detection to reduce hardware complexity and enable platform-agnostic haptic augmentation [11]. In Table 1, these approaches are characterized by high algorithmic complexity but minimal additional hardware, with distal sensing achieved indirectly through image analysis.
@@
Vision-based methods naturally connect to pseudo-haptic and visual-only feedback strategies, where the goal is not to reconstruct forces precisely but to shape the surgeon’s perception of tissue stiffness or danger through visual deformation cues or other modalities. Masui et al. explicitly frame their vision-based force estimation as an attempt to make surgeons’ existing “pseudo-haptic” impressions more explicit and actionable [22]. Trute et al. studied how soft-tissue visual cues in minimal-invasive and robotic surgery can support haptic perception and inform the design of visual augmentations [41]. Broader reviews of haptic technologies for surgical simulation and robotics also stress that pseudo-haptic and multi-modal sensory substitution (e.g., vibrotactile cues) can reduce hardware demands while still improving training and performance [6]. These approaches are highly hardware-agnostic and low-cost but rely heavily on robust computer vision and careful human factor design rather than physically accurate force reconstruction.
@@
Distal force sensing [3, 9, 5] High (H) Low (L) High (H) Distal (D)
Proximal/Joint-space sensing [34, 43, 31] Med (M) Med (M) Robot-side only (R) Proximal (P)
@@
Vision-based force estimation [8, 22, 44, 11] High (H) High (H) Low physical (L) + High data (Hd) Distal (Dind)
Pseudo-haptics / visual cues [22, 41, 6] Low–Med (L–M) Very High (VH) Low (L) Visual distal (V)
Hybrid multi-modal fusion [31, 11, 6, 10] High (H) Med (M) Mixed (Mx) P + D
@@
Taken together, recent literature suggests that there is no single “best” method for robust haptic feedback, instead, different clinical, regulatory, and engineering constraints naturally map to different regions of the design space. For new robotic platforms where instrument redesign is feasible and the cost per-unit is acceptable, instrument-integrated force sensors provide the most direct, distal measurements and the strongest evidence of performance benefits [3, 9, 5]. For retrofitting legacy systems or ensuring cross-platform deployability, proximal sensing and sensorless learning-based models offer a pragmatic compromise between hardware-agnostic deployment and robustness, especially when trained on data that span the expected range of tasks and tissues [34, 43, 31]. Vision-based and pseudo-haptic methods, finally, are attractive when hardware changes are impossible (e.g., closed commercial systems) or when force information is primarily needed for training, monitoring, or safety alerts rather than for continuous kinesthetic feedback [22, 44, 11, 41].
@@
More recently, Impedance Matching Approach (IMA) for force feedback has emerged as an effective strategy for achieving stable and transparent haptic interaction in RAMIS. It operates as a form of sensory substitution, enabling indirect force rendering by bypassing actuators on the leader side and excluding human components from the control loop [45, 26, 35]. Unlike conventional direct force feedback methods, where force error directly commands actuator output, IMA employs closed-loop position control, converting the force error into a corresponding position error that drives the robot [15, 17, 42]. This formulation allows precise regulation of tool–tissue forces while maintaining system stability through an inner position-control loop [39]. Previous studies have validated the IMA framework through simulation and experimental evaluations, demonstrating its potential for stable and accurate haptic rendering [30]. Complementary work has explored integrating tactile object recognition and purposeful haptic exploration to extract object features and improve force perception during teleoperation [33]. Moreover, advanced haptic rendering algorithms have been shown to minimize feedback oscillations and enhance the fidelity of force cues, leading to more realistic and responsive interaction [21]. Building on this, the linear IMA method has recently demonstrated high precision and stability for single-axis force components in surgical teleoperation tasks [37]. We subsequently introduced a preliminary Nonlinear IMA (NIMA) framework [23] that accounts for nonlinear impedance parameters in tool-tissue interactions, demonstrating its feasibility for 3D force rendering with an MAE of 0.03 N and an 85% improvement over linear IMA. However, that initial study did not address tool-tip force isolation from friction at the Remote Center of Motion (RCM), nor did it include coordinate calibration between sensing elements-limitations that the present work addresses through a neural network-based force extraction method and a comprehensive validation protocol.
@@ 1.2 Contributions
This study extends the NIMA framework by addressing critical gaps in tool-tip force isolation and sensor calibration. As illustrated in Fig. 1, the proposed framework continuously identifies nonlinear impedance characteristics in real-time to enable precise rendering of multi-axis contact forces and stable bidirectional communication between the leader and follower modules. The framework introduces a neural network (NN)-based method for extracting true tool-tip forces by isolating frictional components at the RCM, along with a coordinate correspondence calibration procedure between the robot-mounted and tissue-mounted force sensors. The main contributions of this work are:
@@
    A NN-based tool-tip force extraction method that isolates true instrument–tissue interaction forces from frictional artifacts at the RCM.
@@
    A coordinate correspondence calibration procedure enabling accurate transformation between multiple force sensor frames.
@@
    An extended NIMA framework achieving sub-centinewton force rendering accuracy, representing a 95% improvement over linear IMA.
@@
    Comprehensive three-stage experimental validation including coordinate calibration, NN-based force isolation, and NIMA force reconstruction.
  5. 5.
@@
Figure 1: Proposed Nonlinear Impedance Matching Approach (NIMA). The leader module (left) receives motion commands $\mathbf{X}$ from the operator and renders the estimated feedback force $\mathbf{f_{d}}=\mathbf{M}\mathbf{X}$, while the follower module (right) executes $\mathbf{X}$, measures $(\mathbf{f},\mathbf{X})$, and identifies nonlinear impedance parameters $\mathbf{M}$ in real time. This closed-loop structure enables stable, high-fidelity force feedback in robotic laparoscopy [23].
2 Materials and Methods